OpenStudy (anonymous):

slope 2, y-intercept -4

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

(y - y0) = m ( x - x0 ) slope = 2 = m y-intercept -4 x0 = 0 => y0 = -4 y + 4 = 2( x - 0 ) y = 2x -4

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

any Time